# SwarmSearch: Decentralized Search Engine

### v1

@@ V-A Decentralized Search
The search is initiated by a user submitting a query (e.g., “ubuntu download”). The user then solicits matching documents from multiple agents in the network. Once a query is received by an agent, it performs a beam search on its model. The model produces a fixed-sized set of (e.g., 5 or 10) docids that most likely match the query, along with associated logit scores, forming beams in the form $(d_{i},s_{i})$. This result set, subsequently, is sent back to the querying user, who performs a beam search on their own model as well. The scores $s_{i}$ within each set of beams are normalized using the softmax function. Further, all beams are aggregated into a single set of docid candidates, where the scores of equal docids are summed. This procedure approximates the intractable task of fully enumerating the distribution of all possible docid generations. The underlying assumption is that, across an ensemble of similarly trained models, relevant documents will consistently appear among the top candidates from multiple models, even if their exact ranking varies slightly. Consequently, aggregating predictions from multiple models via this ensemble approach yields superior retrieval accuracy compared to relying on a single model. Finally, the top-$k$ docids with the highest scores are selected from this initial retrieval phase to undergo further processing in the re-ranking stage.
@@ V-B Click-based Feedback
When a user selects a document from a search result list, we refer to it as a click. Clicks are central to our system, as they present the sole signal driving our search engine’s training. We interpret clicks as the user deeming a specific document relevant w.r.t. the query. At the same time, we interpret a click as all remaining documents in the search result list being deemed non-relevant w.r.t. the query. These clicklogs serve as feedback to both the retrieval and the ranking model of our search engine, and are further disseminated in the network. The models, therefore, continuously train on more and updated data, enabling them to adapt and improve over time.
@@ VI Dataset
In addition, the training of our ranking model requires the non-clicked documents given a query, i.e., the list of displayed search results given a search query. This data is missing from the dataset and cannot be restored to its original form. To address this limitation, we generate candidate document sets based on the methodology proposed by Guo et al. [61]. Specifically, given each query, we rank the entire set of documents based on the BM25 score of the document titles. BM25 is a standard ranking function used in information retrieval [62, 63]. We then apply a window of size 10 on the ranked list with the clicked document in the center. That is, if the clicked document is ranked at position 20 in the BM25-ordered list, we consider the documents from positions 15–24 the result candidates for the given query. If the position of the clicked document is close to the bounds of the list, we extend the selection accordingly so that we always return 10 candidate documents. In Table II, we show an excerpt of our datasets, consisting of query logs, the clicked document (“Doc ID”), as well as the list of candidates that we compiled.
@@ VII-A Retrieval Accuracy
Figure 2: Top-1 document retrieval accuracy of a single model evaluated on a test set of queries. Each data point represents one epoch of training.
@@ VII-B Spam Prevention
For our experiment, we simulate users connected to 10 randomly sampled users from our dataset; in the following referred to as neighbors. Preliminary experiments revealed converged performance with this number of neighbors. We only consider users with at least three clicklogs. The clicklogs of a user constitute a user’s personal dataset. First, we partition each user’s personal dataset into training, validation, and test sets using a 1:1:1 ratio.
@@
Next, we analyze the effects of adversarial nodes on the ranking performance. To this end, we apply label flipping to $n\in[0,10]$ out of the 10 datasets presented to a user, essentially simulating a poisoning attack [67, 68]. In Figure 3, we show the performance degradation under an increasing share of adversarial nodes $n/10$. We show two modes of operation:
@@ VII-C Resilience Against Sybil Attacks
We simulate a search network of 1000 user nodes, modeled after behavioral patterns in our dataset, where a varying fraction (0–100 %) is controlled by a Sybil adversary. Honest nodes receive reputation scores sampled from our real-world dataset, containing anonymized records of user search activity, including queries, document candidates, and clicks. Each unique User ID is treated as a user node, and reputation scores are computed based on total click activity under the assumption that consistent interaction with search results correlates with meaningful participation in the system. Click counts are normalized to the range $[0,1]$, yielding a continuous reputation distribution that reflects relative contribution levels across the user population. To simulate donations, each round begins with a designated donor node allocating a fixed amount across a subset of peers using the weighted lottery mechanism described in Section V-C. Sybil nodes are assigned low or manipulated reputation scores depending on the scenario, either by sampling from the lower tail of the honest distribution or by injecting dense feedback among themselves. For each configuration, we simulate multiple donation rounds and report the average donation share received by Sybil nodes.

### v2

@@ V-A Decentralized Search
The search is initiated by a user submitting a query (e.g., “ubuntu download”). The user then solicits matching documents from multiple agents in the network. Once a query is received by an agent, it performs a beam search on its model. The model produces a fixed-sized set of (e.g., 5 or 10) docids $d_{i}$ that most likely match the query, along with associated logit scores $s_{i}$, forming beams in the form $(d_{i},s_{i})$. This result set, subsequently, is sent back to the querying user, who performs a beam search on their own model as well. The scores $s_{i}$ within each set of beams are normalized using the softmax function. Further, all beams are aggregated into a single set of docid candidates, where the scores of equal docids are summed. This procedure approximates the intractable task of fully enumerating the distribution of all possible docid generations. The underlying assumption is that, across an ensemble of similarly trained models, relevant documents will consistently appear among the top candidates from multiple models, even if their exact ranking varies slightly. Consequently, aggregating predictions from multiple models via this ensemble approach yields superior retrieval accuracy compared to relying on a single model. Finally, the top-$k$ docids with the highest scores are selected from this initial retrieval phase to undergo further processing in the re-ranking stage.
@@ V-B Click-based Feedback
When a user selects a document from a search result list, we refer to it as a click. Clicks are central to our system, as they present the sole signal driving our search engine’s training. We interpret clicks as the user deeming a specific document relevant w.r.t. the query. Also, we interpret every other document in the search result list as being deemed non-relevant w.r.t. the query. These clicklogs serve as feedback to both the retrieval and the ranking model of our search engine, and are further disseminated in the network. The models, therefore, continuously train on more and updated data, enabling them to adapt and improve over time.
@@ VI Dataset
In addition, the training of our ranking model requires the non-clicked documents given a query, i.e., the list of displayed search results given a search query. This data is missing from the dataset and cannot be restored to its original form. To address this limitation, we generate candidate document sets based on the methodology proposed by Guo et al. [61]. Specifically, given each query, we rank the entire set of documents based on the BM25 score of the document titles. BM25 is a standard keyword-based ranking function used in information retrieval [62, 63]. We then apply a window of size 10 on the ranked list with the clicked document in the center. That is, if the clicked document is ranked at position 20 in the BM25-ordered list, we consider the documents from positions 15–24 the result candidates for the given query. If the position of the clicked document is close to the bounds of the list, we extend the selection accordingly so that we always return 10 candidate documents. In Table II, we show an excerpt of our datasets, consisting of query logs, the clicked document (“Doc ID”), as well as the list of candidates that we compiled.
@@ VII-A Retrieval Accuracy
Figure 2: Top-1 document retrieval accuracy of a single model evaluated on a test set of queries. Each data point represents one epoch of training. The x-axis is truncated where convergence was observed across all lines.
@@ VII-B Spam Prevention
For our experiment, we simulate users connected to 10 randomly sampled users from our dataset; in the following referred to as neighbors. Preliminary experiments revealed converged performance with this number of neighbors. We only consider users with at least three clicklogs. The clicklogs of a user constitute a user’s personal dataset. First, we partition each user’s personal dataset into training, validation, and test sets using a 1:1:1 ratio. This split ratio was chosen to maximize the number of users meeting the minimum data threshold.
@@
Next, we analyze the effects of adversarial nodes on the ranking performance. To this end, we apply label flipping to $n\in[0,10]$ out of the 10 datasets presented to a user, essentially simulating a poisoning attack [67, 68]. In Figure 3, we show the performance degradation under an increasing share of adversarial nodes $n/10$. We show three modes of operation:
@@
  • Oracle: For reference, we report an oracle baseline that excludes any poisoned datasets.
@@ VII-C Resilience Against Sybil Attacks
We simulate a search network of 1000 user nodes, modeled after behavioral patterns in our dataset, where a varying fraction (0–100 %) is controlled by a Sybil adversary. Honest nodes receive reputation scores sampled from our real-world dataset, containing anonymized records of user search activity, including queries, document candidates, and clicks. Each unique “User ID” is treated as a user node, and reputation scores are computed based on total click activity under the assumption that consistent interaction with search results correlates with meaningful participation in the system. Click counts are normalized to the range $[0,1]$, yielding a continuous reputation distribution that reflects relative contribution levels across the user population. To simulate donations, each round begins with a designated donor node allocating a fixed amount across a subset of peers using the weighted lottery-based mechanism described in Section V-C. Sybil nodes are assigned low or manipulated reputation scores depending on the scenario, either by sampling from the lower tail of the honest distribution or by injecting dense feedback among themselves. For each configuration, we simulate multiple donation rounds and report the average donation share received by Sybil nodes.